(7) A man borrows a loan of $20,000 to purchase a car at annual rate of interest of 6%. he will pay back the loan through monthly installment over 5 years with the first installment to be made three months after the release of the loan. what is the monthly installment he needs to pay?
Answer #1

It is assumed that the loan account will be closed over 5 years, so repayment should be in 57 installments (ie. 5*12 - 3).

Formula : Installment = P (i/c) (1+i/c)^nc / (1+i/c)^nc - 1

Installment = 20000* (0.06/12) * (1+0.06/12)^57 / (1+0.06/12)^57 - 1

Installment = 20000* 0.005 * (1+0.005)^57 / (1+0.005)^57 - 1

Installment = 100 * 1.328818 / 1.328818 - 1

Installment = 132.8818 / 0.328818 = $404.12